I need help with this problem.
3x+4/5 = x-3/2 please send help to anevarez@psinc3.com.
Thanks,
Al

Answer by renevencer22(21)


3*x+ (4/5) = x-(3/2)
mult thru by 10: 30*x + 4(2) = 10*x - 3*5
20*x = -23
x = -23/20 = -1.15
To check:
3*(-1.15) + (4/5) = -1.15 -(3/2)
-2.65 = -2.65